### v1

# Spatial Knowledge Graph-Guided Multimodal Synthesis

Abstract
Recent advances in multimodal large language models (MLLMs) have significantly enhanced their capabilities; however, their spatial perception abilities remain a notable limitation. To address this challenge, multimodal data synthesis offers a promising solution. Yet, ensuring that synthesized data adhere to spatial common sense is a non-trivial task. In this work, we introduce SKG2Data, a novel multimodal synthesis approach guided by spatial knowledge graphs, grounded in the concept of knowledge-to-data generation. SKG2Data automatically constructs a Spatial Knowledge Graph (SKG) to emulate human-like perception of spatial directions and distances, which is subsequently utilized to guide multimodal data synthesis. Extensive experiments demonstrate that data synthesized from diverse types of spatial knowledge, including direction and distance, not only enhance the spatial perception and reasoning abilities of MLLMs but also exhibit strong generalization capabilities. We hope that the idea of knowledge-based data synthesis can advance the development of spatial intelligence.
Spatial Knowledge Graph-Guided Multimodal Synthesis
Yida Xue1,  Zhen Bi1,2,  Jinnan Yang2,  Jungang Lou2, Huajun Chen1,  Ningyu Zhang1thanks:  Corresponding Author. 1Zhejiang University  2Huzhou University {xueyida, bizhen_zju, huajunsir, zhangningyu}@zju.edu.cn
1 Introduction
Despite significant advancements in Multimodal Large Language Models (MLLMs) Zhao et al. (2023b); Yin et al. (2023) for visual processing tasks, a critical limitation persists in their ability to comprehend spatial relationships Prasad et al. (2023); Liu et al. (2023b); Du et al. (2024); Lei et al. (2024); Tong et al. (2024); Shiri et al. (2024); Zheng et al. (2024). Tong et al. (2024) observe that MLLMs often operate with “eyes wide shut”, revealing fundamental deficiencies in visual-spatial comprehension even in state-of-the-art models. This limitation underscores a significant gap between MLLMs and human capabilities, as humans inherently excel in tasks requiring spatial intelligence.
Figure 1: The Spatial Knowledge Graph (SKG) serves as a structured representation of object attributes and spatial relationships. Guided by the SKG, SKG2Data generates images and multimodal data, grounded in the concept of knowledge-to-data generation.
To address this challenge, multimodal data synthesis offers a promising solution, requiring the generation of large-scale, spatially coherent multimodal data that accurately adheres to real-world constraints. A straightforward approach involves developing computational models (or world models) capable of generating continuous synthetic data Du et al. (2024); Ray et al. (2024); Li et al. (2025). However, this method is inefficient due to redundant data across scenarios and demands frequent model updates when environments change, leading to significant maintenance costs. Notably, human spatial intelligence stems from the brain’s acquisition of spatial knowledge, which enables robust generalization through the application of cognitive frameworks Hodges et al. (1999); Herweg and Kahana (2018); Peer et al. (2021). This insight naturally leads to a new paradigm for multimodal data synthesis: modeling spatial knowledge to guide the targeted generation of informative data.
Inspired by this insight, as illustrated in Figure 1, we introduce SKG2Data, an innovative framework specifically designed to generate multimodal synthetic data for spatial understanding. The core idea is to leverage a specialized knowledge graph (KG) as a spatial knowledge representation to guide the generation of multimodal data. Specifically, we construct a Spatial KG (SKG) that simulates human-like understanding of spatial knowledge by characterizing scenes and modeling spatial relationships. SKG2Data utilizes the SKG, which encodes object attributes and their spatial relationships, to generate corresponding images and textual data.
We automatically construct a multimodal instruction dataset for MLLMs based on the SKG, comprising 974 nodes and 4,731 triplets that encode spatial relationships. To ensure a fair comparison of results and prevent data contamination, we also use SKG2Data to synthesize an evaluation dataset, referred to as SKG2Data-Holdout, which includes 120 images and 566 single-choice questions. We fine-tune the LLaVA-1.6-Vicuna-7B Liu et al. (2024a) and Llama-3.2-Vision-11B Meta AI (2024) models using the synthesized data. The fine-tuned models show significant performance improvements on SKG2Data-Holdout and other spatial understanding benchmarks, while maintaining their general abilities. Additionally, we carry out comprehensive experiments to examine the effect of different spatial relation types, dataset sizes, and object densities on model performance.
Figure 2: A comprehensive overview of the our framework. Our framework consists of two core modules: Spatial KG Generation and Multimodal Data Synthesis. The Spatial KG Generation module generates an intermediate representation, the Spatial KG, which guides the synthesis of multimodal data. The Multimodal Data Synthesis module is tasked with generating Image Data and their corresponding Textual Data.
2 Preliminaries
Spatial Knowledge Graph.
In recent years, KGs have been widely applied across various domains, such as FACTKG Kim et al. (2023) for fact verification. Unlike traditional KGs, we introduce the Spatial KG (SKG), a specialized KG designed to encode spatial knowledge and guide multimodal data synthesis for spatial tasks. The SKG is defined as a KG where nodes represent objects and edges represent spatial relationships through triplets. Formally, SKG can be represented as:
$G=(E,T)$ (1)
where $E$ denotes the set of nodes (objects) and $T$ denotes the set of triplets describing spatial relations. These relationships may include directional relationships (e.g., “to the left of” and “to the right of”), distance relationships (e.g., “close to” and “far away”), or a combination of both. SKG represents a collection of objects and their spatial relationships within a specific scene.
In contrast to Scene Graphs Zhu et al. (2017, 2022), which provide a general representation of scenes, the triplets in an SKG are specifically designed to emphasize spatial relations, with nodes $E$ containing detailed descriptions of objects. This structured representation enables the generation of complex and detailed multimodal synthesis data that emphasize spatial relationships.
Multimodal Data Synthesis.
The task of multimodal data synthesis involves generating high-quality input-output pairs related to images using a model $f$. For a given image $i$ (either pre-existing or synthesized), the task uses $i$, its description $d$, or a combination of both as inputs to $f$. The model $f$ then synthesizes a set of input $Q$ and their corresponding output $A$.
We focus on generating multimodal data, with an emphasis on capturing spatial relationships between objects. Our methodology involves the simultaneous creation of images and their corresponding question-answer pairs through a three-component framework: a multimodal generative model $f_{1}$, a text-to-image generative model $f_{2}$, and a SKG $G$. Inspired by LLM-grounded Diffusion (LMD) Lian et al. (2024), we use $f_{1}$ to generate bounding boxes $B$ and description $d$ under the guidance of $G$. These bounding boxes are then integrated into the layout-diffusion model $f_{2}$ to synthesize images. After image synthesis, $G$ and $B$ further assist $f_{1}$ in generating multiple question-answer pairs $(Q,A)$ relevant to the synthesized image $i$.
$\begin{gathered}(B,d)=f_{1}(G)\\ i=f_{2}(B,d,G)\\ (Q,A)=f_{1}(B,G,i)\end{gathered}$ (2)
For implementation, we use GPT-4o and GLIGEN Li et al. (2023) version of LMD as $f_{1}$ and $f_{2}$, respectively.
3 Spatial Knowledge Graph-Guided Data Synthesis
In this section, we introduce SKG2Data, a novel approach that takes advantage of a SKG that incorporates spatial positional relationships to guide the synthesis of multimodal data. As shown in Figure 2, SKG2Data consists of two core modules: Spatial KG Generation and Multimodal Data Synthesis.
3.1 Spatial Knowledge Graph Generation
Scenes and Objects Generation.
The real world encompasses a wide variety of scenes, each containing distinct sets of objects. To simulate the distribution of scenes and objects in real-world images, we leverage GPT-4o to generate a series of scenes and the objects likely to appear within them. Starting with a few scenes as few-shot examples, we use GPT-4o to generate an additional 160 safe, diverse, and realistic scenes, significantly enhancing the diversity of our dataset. For each scene, GPT-4o is further employed to produce a list of plausible objects that could exist in the given context. To improve object generation, particularly for uncommon scenes, we incorporate Wikipedia documents as external knowledge to augment GPT-4o’s understanding. This approach ensures that the synthesized objects exhibit distributions closely aligned with real-world images, effectively preventing the generation of incongruous or improbable objects within the specified scenes. Formally, we first employ $f_{1}$ to generate $m$ general scenes, denoted as $\{s_{1},s_{2},\dots,s_{m}\}$. For each scene $s_{i}$, we then generate $n$ objects that are likely to appear in that scene, expressed as:
$\{o_{i1},o_{i2},\dots,o_{in}\}=f_{1}(s_{i})$ (3)
Figure 3: Distribution of top 15 objects and spatial relationships. There are a total of 974 objects and 95 relationships.
Spatial Knowledge Graph Construction.
Due to the diverse rules and knowledge governing spatial relationships between objects in the real world, we employ GPT-4o to generate positional relationships that align with real-world spatial cognition. Additionally, we introduce the SKG as an intermediary representation to facilitate the generation of high-quality multimodal data for spatial understanding tasks. The SKG is constructed using GPT-4o’s capabilities to synthesize entities and their spatial relationship triplets.
In the initial phase, $f_{1}$ is used to select a subset $\{o_{ij_{1}},o_{ij_{2}},\ldots,o_{ij_{k}}\}$ of $k$ candidate objects from the original set $\{o_{i1},o_{i2},\ldots,o_{in}\}$. These selected objects are utilized to construct a specific SKG. The subset $C_{k}$ is formally represented as:
$\begin{gathered}C_{k}=\{o_{ij_{1}},o_{ij_{2}},\ldots,o_{ij_{k}}\},\\ C_{k}\subseteq\{o_{i1},o_{i2},\ldots,o_{in}\},\\ |C_{k}|=k\end{gathered}$ (4)
Next, each object in $C_{k}$ undergoes an attribute enrichment process via $f_{1}$, which adds quantities and attributes such as color, orientation, and material. This results in a set of $l$ entities, denoted as $E_{l}$, with rich descriptive attributes. In $E_{l}$, each entity represents a single object with a quantity of one and clearly defined attributes. If objects within the same category have a quantity greater than one, they are divided into multiple entities. For example, a selected object “Balloon” with a higher quantity could be expanded into two distinct entities which reflect the addition of specific attributes by GPT-4o: “A blue balloon” and “A yellow balloon”. This approach ensures precise control over the number of entities, maintaining consistency with the objects in the subsequently generated images.
$\begin{gathered}E_{l}=\{e_{ij_{1}},e_{ij_{2}},\ldots,e_{ij_{l}}\}=f_{1}(C_{k}),\\ |E_{l}|=l\geq k\end{gathered}$ (5)
The entities in $E_{l}$ serve as the nodes of the SKG $G$. The subsequent task involves synthesizing spatial relationships between pairs of entities in $E_{l}$. This is achieved using $f_{1}$, which generates a set of relationship triplets $T_{l}$:
$\begin{gathered}T_{l}=\{(e_{ij_{a}},r,e_{ij_{b}})\mid e_{ij_{a}},e_{ij_{b}}\in E_{l},\\ a\neq b,r=f_{1}(e_{ij_{a}},e_{ij_{b}})\}\end{gathered}$ (6)
Each triplet $(e_{ij_{a}},r,e_{ij_{b}})$ represents a unique pair of entities $e_{ij_{a}}$ and $e_{ij_{b}}$, along with their spatial relationships $r$ which could involve directional relationships, distance relationships, or a combination of both. The collection of these triplets $T_{l}$ comprehensively captures all pairwise spatial interactions within $E_{l}$. By combining the nodes $E_{l}$ and the triplet relationships $T_{l}$, we construct the $G$.
3.2 Multimodal Data Synthesis
Image Data Generation.
End-to-end text-to-image generative models, such as Stable Diffusion Rombach et al. (2021) and Dall-E Betker et al. (2023), often encounter challenges in accurately controlling object positions and quantities. In contrast, layout-based diffusion models enable precise control over object placement and quantities, ensuring that the generated images closely conform to SKG. We employ GPT-4o to generate bounding boxes and captions under the guidance of SKG, and seamlessly incorporate these outputs into a GLIGEN-based model. Recognizing that diffusion models cannot guarantee perfection in every generated image, we generate multiple images for each SKG by varying the random seed.
Specifically, we input the intermediate representations $E_{l}$ and $T_{l}$ derived from $G$ directly into $f$ as guidance. Additionally, we leverage few-shot examples to instruct $f_{1}$ on how to generate appropriate bounding boxes $B_{l}$.
$\begin{gathered}B_{l}=\{b_{ij_{1}},b_{ij_{2}},\ldots,b_{ij_{l}}\}=f_{1}(E_{l},T_{l}),\\ \end{gathered}$ (7)
In order to synthesize data on whether an object exists in the image for subsequent tasks, we randomly select an object from other objects in current scene as a non-existent object $o_{neg}$, ensuring that this object is excluded from the synthesized image. The elements $E_{l}$, $T_{l}$, and $s_{i}$, along with the optional non-existent object $o_{neg}$, are input to $f_{1}$ to produce the caption $d$.
$\begin{gathered}o_{neg}\in\{o\mid o\notin C_{k}\}\cup\{\emptyset\},\\ d=f_{1}(E_{l},T_{l},s_{i},o_{neg})\end{gathered}$ (8)
Finally, we input $B_{l}$ and $d$ into $f_{2}$ to generate the final image $i$.
$\begin{gathered}i=f_{2}(d,B_{k})\end{gathered}$ (9)
Image Data Filtering.
During the image generation stage, a significant challenge is the potential for discrepancies, such as hallucinations, which can occur during the transformation of the SKG into bounding boxes and captions. Additionally, the image generation model may introduce its own inaccuracies. These issues can result in inconsistencies between the generated image and the original SKG. To address this challenge, we conduct an image validation process using GPT-4o as a image filtering agent. This process is aimed at evaluating and ensuring that the generated image aligns with the SKG. Formally, we utilize $f_{1}$ to assess the alignment of an image $i$ with the expected entities $E_{l}$ and textual descriptions $T_{l}$. If $f_{1}$ returns True, the image $i$ is preserved; otherwise, it is discarded. This process is represented by the following equation:
$f_{1}(i,E_{l},T_{l})=\begin{cases}\text{True},&\text{if }i\text{ aligns with}\\ &\text{both }E_{l}\text{ and }T_{l},\\ \text{False},&\text{otherwise.}\end{cases}$ (10)
After filtering, the dataset contains a total of 974 objects and 95 spatial relationships. The distribution of the top 15 objects and spatial relationships in the filtered images is illustrated in Figure 3.
Dataset LLaVA-1.5 LLaVA-1.6 LLaVA-1.6 (w/ SKG2Data) Llama-3.2-vision Llama-3.2-vision (w/ SKG2Data) KG2Data-Holdout 62.2 68.6 70.1 (+1.5) 73.3 74.7 (+1.4) COCO-Spatial 52.4 75.4 79.3 (+3.9) 45.9 59.8 (+13.9) MMVP 24.7 32.0 36.7 (+4.7) 29.3 30.7 (+1.4) MMStar 33.1 37.6 36.7 (-0.9) 49.8 48.1 (-1.7) HallusionBench 27.6 27.6 27.2 (-0.4) 40.3 45.1 (+4.8) Average 40.0 48.2 50.0 (+1.8) 47.7 51.7 (+4.0)
Table 1: Comparison with SoTA methods on five benchmarks.
Textual Data Generation.
The training data for textual data is primarily composed of open-ended question-answer pairs, while the evaluation data predominantly consists of single-choice questions. To enhance GPT-4o’s understanding of object attributes and spatial relationships in images, we leverage the SKG to guide the generation of question-answer pairs. This approach not only improves the accuracy of the data generation process but also reduces the occurrence of hallucinations. Furthermore, it enables the creation of more diverse and complex questions tailored for spatial understanding tasks. These questions are categorized into two types based on their primary focus: Entity-Based Data and Relation-Based Data.
[…]
$\begin{gathered}x=f_{1}(i,B_{l},E_{l},T_{l}),\\ y=f_{2}(i,B_{l},E_{l},T_{l},x)\end{gathered}$ (11)
Textual Data Filtering.
For each generated textual data, we employ GPT-4o as a text filtering agent to verify its accuracy. To avoid excessive reliance on the information provided in the SKG, we withhold SKG during the textual data verification phase. Instead, GPT-4o relies solely on the image to verify the correctness of the provided reference answers. The evaluation is performed using $f_{1}$, which determines the correctness of the answer $y$ for the question $x$ based on the image $i$. If $f_{1}$ returns True, the textual data $(x,y)$ is preserved:
$f_{1}(i,x,y)=\begin{cases}\text{True}&\text{if }y\text{ is correct},\\ \text{False}&\text{otherwise.}\end{cases}$ (12)
Examples of the final training data generated are illustrated in Figure 7 (Appendix A).
4 Experiments
4.1 Experimental Settings
Baselines.
We conduct experiments on three open-source MLLMs: LLaVA-1.5-7B Liu et al. (2023a), LLaVA-1.6-7B Liu et al. (2024a), and Llama-3.2-Vision-11B Meta AI (2024). Both LLaVA-1.5 and LLaVA-1.6 use the Vicuna-7B Chiang et al. (2023) LLM as their LLM backbone.
Benchmarks.
For a comprehensive evaluation, we examine all models across two distinct task categories. The first category encompasses benchmarks that are either directly or partially related to spatial understanding. This includes the publicly available benchmark MMVP and COCO-Spatial Kamath et al. (2023). For COCO-Spatial, we utilize a subset named Two-obj that focuses on the positional relationships of two objects. We then convert the original positive and negative sample pairs into single-choice questions, which are more conducive to MLLMs testing. We leave a portion of the synthesized image to generate test questions called SKG2Data-Holdout, which are also used for evaluation. The second category involves general capability testing benchmarks for MLLMs, including MMStar and HallusionBench.
Implementation Details.
We fine-tune both LLaVA-v1.6 and Llama-3.2-Vision using training data. These models consist of a vision encoder, a projector, and an LLM backbone. During training, we tune the LLM backbone with LoRA modules  Hu et al. (2021), keeping the projector and vision encoder parameters fixed. The detailed settings are shown in Table 2 (Appendix A).
4.2 Main Results
Table 1 presents a comprehensive comparison of multiple models evaluated across five benchmarks, highlighting the strengths of SKG2Data.
Spatial Understanding Results.
Fine-tuned LLaVA-1.6 and fine-tuned Llama-3.2-vision demonstrate marked improvements across several challenging benchmarks for spatial tasks, highlighting the efficacy of our methods. On the KG2Data-Holdout benchmark, Fine-tuned LLaVA-1.6 achieves a score of 70.1(+1.5)↑, while Fine-tuned Llama-3.2-vision reaches 74.7(+1.4)↑, indicating robust performance enhancements. For the COCO-Spatial benchmark, Fine-tuned LLaVA-1.6 outperforms with a score of 79.3(+3.9)↑, and Fine-tuned Llama-3.2-vision impressively achieves 59.8(+13.9)↑, suggesting significant spatial understanding advancements. In the MMVP benchmark, Fine-tuned LLaVA-1.6 shows notable improvement with a score of 36.7(+4.7)↑, while Fine-tuned Llama-3.2-vision also makes gains, reaching 30.7(+1.4)↑. In summary, our experimental results indicate that synthesized training data can improve the spatial understanding ability of MLLMs.
General Visual Understanding Results.
We present the results of our models across several general visual understanding benchmarks: MMStar, MMVet, and HallusionBench. In the MMStar benchmark, fine-tuned LLaVA-1.6 slightly decreases to 36.7, reflecting a minor decline of 0.9↓, while fine-tuned Llama-3.2-vision shows a slight decrease to 48.1, marking a reduction of 1.7↓. Interestingly, in the HallusionBench benchmark, fine-tuned LLaVA-1.6 experiences a negligible decrease to 27.2, showing a marginal drop of 0.4↓, while fine-tuned Llama-3.2-vision demonstrates a remarkable improvement, advancing to 45.1(+4.8)↑, indicating enhanced performance in hallucination task. Overall, our results demonstrate that training LLaVA-1.6 and Llama-3.2-vision with synthesized data does not significantly degrade their general visual understanding abilities.
Average Results.
The fine-tuned LLaVA-1.6 model achieves an average score of 50.0 (+1.8), while the fine-tuned Llama-3.2-vision model reaches 51.7 (+4.0). These improvements confirm that SKG2Data effectively enhances spatial reasoning while maintaining performance in general visual understanding tasks.
Figure 4: The result of removing specific positional relationship data with the same amount of data.
Figure 5: The influence of the number of objects and the quantity of data on experimental results.
4.3 Ablation and Analysis
To evaluate the impact of the synthesized data generated by SKG2Data on model performance, we conducted a comprehensive ablation study using LLaVA-1.6. This study focuses on three key aspects: (1) the influence of different relationship types, (2) the effect of varying data quantities, and (3) the impact of the number of objects.
For the first aspect, GPT-4 is utilized to categorize the positional relationships in the training data into two types: distance relationships and directional relationships. The models are then trained separately on each type. To minimize the influence of data quantity on the experimental results, we sample 2,000 data from the training dataset after eliminating one type of positional relationship. The detailed experimental results are depicted in Figure 4. For the second aspect, we randomly sample 2k and 5k from the 15k training data. For the third aspect, we regulate the data volume to 5k to examine the influence of the number of objects on the model performance within SKG. Setting the limit to 3 objects, we categorize the numbers into two groups: (1) counts that are greater than or equal to 3 and (2) those that are less than 3. The results of second and third aspects are shown in Figure 5. The experimental results shown in Figure 4 guide us towards Finding 1 and Finding 2. Additionally, Finding 3 is derived from the experimental results observed in Figure 5.
Figure 6: Case analysis: On spatially related datasets, the trained model enhances spatial understanding abilities.
Finding 1: Directional knowledge plays a crucial role in enhancing the MLLM’s spatial perception capabilities.
We observe that training data containing directional relationships leads to more significant performance improvements compared to data with distance relationships across three spatially correlated datasets. This suggests that MLLMs may be more sensitive to directional knowledge and prioritize spatial perception and understanding through directional relationships. A potential explanation is that directional knowledge, compared to distance knowledge, exhibits greater variability, providing stronger spatial signals that are more effective in enhancing model performance.
Finding 2: Data synthesized from two types of spatial knowledge, including direction and distance, exhibit generalization ability.
It can be observed that the use of two types of spatial relationship data can enhance performance across three datasets. Significantly, the COCO-Spatial dataset primarily targets directional understanding, and employing data that solely comprises distance relationships for training models can also enhance performance on this dataset. This improvement could possibly be attributed to the varied spatial correlations present in the data synthesized through the spatial knowledge in SKG, and the generalization of an array of positional relationships. This phenomenon underscores the superiority of knowledge-based data synthesis over previous methods.
Finding 3: Increasing the number of objects improves the performance of MLLMs in visual detail tasks
Our experiments demonstrate that synthesizing data with varying quantities of objects improves spatial understanding, with greater gains observed as the number of objects increases in MMVP benchmark. Notably, training with more objects significantly enhances fine-grained recognition capabilities, as reflected in improved performance on the detail-oriented MMVP benchmark. We attribute these improvements to the inherent complexity of multi-object scenes, which expose the model to richer spatial relationships and constraints during learning. By leveraging the structured spatial knowledge embedded in SKG, our proposed SKG2Data systematically generates multimodal training data with diverse and spatially plausible object arrangements, enabling more effective learning of visual and spatial reasoning.
Discussion on Knowledge-to-Data Synthesis.
It can be seen that with more data, the performance of the trained model can be further improved. At the same time, it can also be found that high-quality data is important, and a small amount of high-quality training data can significantly improve the performance of models in specific fields, which is in line with the opinion of “less is more” Zhou et al. (2023); Ye et al. (2025). As shown in Figure 6, the trained model exhibits improved accuracy on spatially related questions. The core idea of our work is to synthesize data through automated modeling knowledge to ensure the rationality of the data itself. The data generated through spatial knowledge is consistent with human spatial cognition, ensuring the relevance and reliability of the data. In addition, to synthesize additional data, we can either randomly sample more data from the SKG or expand the scenes to further enrich the SKG. Compared with the method of synthesizing multimodal data using code, simulators, etc., it is more efficient and has better generalization ability.
5 Related Work
Multi-Modal Large Language Model.
The remarkable success of large language models (LLMs) Brown et al. (2020); Achiam et al. (2023); Touvron et al. (2023) has driven the development of MLLMs Ye et al. (2023); Liu et al. (2024a); Chen et al. (2023); Meta AI (2024); Gemini Team (2024); Bai et al. (2023); OpenAI (2023, 2024). By incorporating visual backbones, these MLLMs demonstrate strong performance across a wide range of visual-language tasks, significantly improving the processing and understanding of both text and visual information. At the same time, a variety of benchmarks Yu et al. (2023); Guan et al. (2024); Chen et al. (2024) have been developed to evaluate the capabilities of large multimodal models. Additionally, with increasing attention on spatial understanding in MLLMs, specialized benchmarks Kamath et al. (2023); Du et al. (2024); Tong et al. (2024) have been introduced. These benchmarks are specifically designed to tackle the challenges of evaluating spatial comprehension in MLLMs.
Synthetic Data Generation.
Synthetic data has garnered significant attention as a novel solution to the challenges of obtaining large, diverse, and high-quality datasets Chang et al. (2024); Liu et al. (2024b); Long et al. (2024); Kim et al. (2024). Many existing studies primarily rely on pre-existing images Wu et al. (2024), occasionally incorporating additional image information such as captions, bounding boxes, or OCR input Liu et al. (2023a); Zhao et al. (2023a); Zhang et al. (2023). However, another area of research transcends generating textual instruction data and focuses on creating high-quality images through various synthesis techniques rather than relying solely on existing image datasets. For example, REACHQA He et al. (2024) and Multimodal Self-Struct Zhang et al. (2024) use code to precisely synthesize chart images. Additionally, EmbSpatial utilizes simulators to generate images, while SynCLR Tian et al. (2024) and VisMin Awal et al. (2024) employ diffusion models to produce new images. Furthermore, due to the tendency of LLMs to exhibit biases and follow a long-tailed distribution due to their lack of specialization in specific domains Navigli et al. (2023); Fei et al. (2023), there is a growing body of research Feng et al. (2023); Xu et al. (2024); Wang et al. (2024); Yang et al. (2025); Ovadia et al. (2025) that employs knowledge enhancement techniques, such as KG, to generate higher-quality data.
6 Conclusion
In this work, we propose SKG2Data, a novel multimodal synthetic data generation method that leverages Spatial Knowledge Graph to synthesize images and question-answer pairs, addressing the lack of high-quality data for spatial understanding tasks. Using SKG2Data, we create a multimodal instruction dataset and a benchmark, SKG2Data-Holdout, to enhance and evaluate the spatial understanding capabilities of MLLMs. Extensive experiments demonstrate the effectiveness of SKG2Data and analyze the impact of spatial relationship types, dataset scales, and object densities.
Limitations
Noise in synthetic data.
Despite the implementation of automatic filtering, the synthesized data may still contain noise. This noise arises from potential mismatches between images and text in diffusion models, as well as the inherent possibility of hallucinations in LLMs.
Difficult to synthesize data with a large number of objects.
Given the limitations of current diffusion models, generated images can sometimes show problems like distortion and inconsistencies in object type, attributes, quantity, or spatial arrangement. The challenge of creating realistic images grows substantially as the number of objects in the scene increases. In some real-world situations where scenes are packed with many objects, our method may have difficulty accurately depicting these complex scenes.
Appendix A Appendix
Hyperparameter Setting
Fine-tuning method LoRA
Gobal Batch Size 64
[…]
Learning Rate (LR) $51e-$ 5
LoRA $r$ 16
LoRA $\alpha$ 32
LoRA Dropout 0.05
Epoch 2
Mixed precision Enabled
Table 2: Detailed experimental settings.
Figure 7: Training data cases.
Prompt for "Scenes Generation":
You are an AI assistant specializing in imagining the scenes depicted in photos. Your task is to envision a variety of scenes that may appear in different photos. Your response needs to meet the following requirements: 1. Ensure that every scene is safe, accessible to most people, and common in real and daily life scenes. 2. The scenes in the photos should be diverse and non repetitive. Each generated scene typically depicts a facility (usually requiring interaction with multiple objects such as a table or furniture) or a room, or a place or area. Please directly output the generated scenes in the following list format: <In context learning> Now, please generate 160 different scenes. <image>
Prompt for "Objects Generation":
<Wikipedia documents> You are an AI assistant specializing in identifying objects likely to appear in a given scene. Your task is to list as many objects as possible that may be found in that scene. Your response needs to meet the following requirements: 1. Ensure that each object is based on common situations in reality and daily life. Ensure that objects are relatively simple and easy to draw, and once you see their images, you can determine what objects they are. Ensure that the object is safe and harmless. 2. The objects provided should not be duplicated or too similar. Try to describe an object with one word. If there are words with multiple meanings, it is necessary to ensure that they are distinguishable, such as "animal mouse" and "computer mouse". 3. Each scene needs to list as many objects as possible. If there may be people on site, please add ’person’. Here I will give you some examples for reference and use the following template in the format of a list for the response: <In context learning> <image>
Prompt for "Spatial KG Generation (Nodes)":
You are an AI assistant specializing in describing objects in a given scene. Your task is to first select a subset of candidate objects, and then describe the attributes of the selected objects. Your response should meet the following requirements: 1. Select 2 to 4 suitable objects from the given objects to form a possible combination. Note that for some candidate objects that are suitable for repeated appearance, they can be selected repeatedly. Please prioritize selecting more common objects. Do not choose objects that are difficult to represent or recognize in the image. 2. Provide further attribute descriptions for the selected objects, including color, orientation, etc., and describe countable objects as singular. Be careful not to describe objects that do not exist in the real world or are unsuitable for the current scene. The description of each object should be concise, without being overly complex and abstract, and avoid including details that cannot be visualized in the image. For some objects that are difficult to add descriptions, additional descriptions can be omitted. 3. The selected combination of objects can be repeated, but the described attributes need to be different. 4. If there are words with multiple meanings, it is necessary to ensure that they are distinguishable and you need to determine which one is more likely based on the scene and other objects, such as ’animal mouse’ and ’computer mouse’. 5. You need to provide 25 to 50 possible combinations based on the number of objects given. Here I will give you some demonstrations for reference and use the following list template for the response: <In context learning> <image>
Prompt for "Spatial KG Generation (Triplets)":
You are an AI assistant specializing in describing the positional relationship between objects in a scene. Your task is to describe the positional relationship of the given objects. Your response needs to meet the following requirements: 1. Be careful not to describe positional relationships that do not exist in the real world. When describing the positional relationship between two objects, do not involve other objects. At most one positional relationship triplet is generated between every two objects. 2. There may be two types of positional relationships between two objects: directional relationships and distance relationships. Directional relationships include above, below, left, right, in front of, behind, etc. Distance relationships include close, far, near, etc. Separate two positional relationships with commas when generating them simultaneously. For situations with a small number of objects, prioritize generating directional relationships. Here I will give you some demonstrations for reference and use the following template in the format of a list for the response: <In context learning> <image>
Table 3: The prompt examples that we use in the first module.
Prompt for "Image Data Generation":
You are an intelligent bounding box generator. I will provide you with the scene, objects and their spatial relationships in a photo. Your task is to generate the bounding boxes for the given objects, along with a background prompt describing the scene and a prompt describing the photo. The images are of size 512x512. The top-left corner has coordinate [0, 0]. The bottom-right corner has coordinnate [512, 512]. The bounding boxes should not overlap or go beyond the image boundaries. Each bounding box should be in the format of (object name, [top-left x coordinate, top-left y coordinate, box width, box height]) and should not include more than one object. Do not put objects that are already provided in the bounding boxes into the background prompt. Do not include non-existing or excluded objects in the background prompt. Use "A realistic scene" as the background prompt if no background is given in the prompt. There may be multiple spatial relationships between objects, and special attention should be paid to using bounding boxes to represent these spatial relationshipss. Please refer to the example below for the desired format. <In context learning> <image>
Prompt for "Image Data Filtering and Textual Data Generation":
You are an AI assistant specializing in evaluating the quality of a given image and generating question and answer pairs related to the image. Your task is to first determine whether the given image should be saved based on its quality and whether it matches the descriptions, and then continue generating question and answer pairs if it is determined to be saved. Descriptions of the image will provide information about the objects present in the image, their quantity, their attributes, their positional relationships, their bounding boxes (object name, [top-left x coordinate, top-left y coordinate, box width, box height]), and objects that are absent from the image. You need to follow the steps below to complete the task: Step 1. You need to ensure the clarity, quality, and authenticity of the image. You need to verify if there are any unrealistic or unreasonable parts in this image, such as clearly unrealistic looking humans or animals, and if there are any indistinguishable objects. Step 2. You need to ensure that the image contains the specified objects and does not contain any excluded objects. Specified objects are given in the form of dict, and excluded objects are presented in a list format. Please note that the image should accurately reflect the attributes of the specified object, such as Orientation and Direction, State and Condition, Color and Appearance, Structural and Physical Characteristics, Viewpoint and Perspective. Step 3. You need to ensure that the objects in the image have a positional relationship with the matching descriptions. The positional relationships are given in the form of a list, and the objects correspond to the dict. Step 4. If the image matches step 1 and basically matches steps 2 and 3, please answer ’Yes’ and proceed to the next steps to generate question answer pairs. Otherwise, please answer ’No’ and do not proceed to the next steps. Step 5. You need to generate as many question answer pairs as possible based on the correct parts of these descriptions and information that can be confidently determined in the given image, while ensuring quality and diversity. For Step 5, the following requirements must be met: 1. Please note not to mention any information related to the image descriptions in your question-answer pairs, as the image description will not be provided in future tasks using these question-answer pairs. 4. Use the template in the format of a JSON list to generate question and answer pairs without any other information. Below are reference responses for saving and not saving this image: <In context learning> <image>
Prompt for "Textual Data Filtering":
You are an AI assistant specializing in reviewing question-answer pairs about a given image. Your task is to find the correct question-answer pairs based on the given image. Your response needs to meet the following requirements: 1. You need to determine whether the answers to the questions are correct based on the given images, and save the correct question-answer pairs. 2. Use a JSON list format template with the same format as the input to generate correct question-answer pairs without generating any additional information. Here is the question and answer about the given image: <In context learning> <image>
Table 4: The prompt examples that we use in the second module.

### v2

@@
 Yet, ensuring that synthesized data adhere to spatial common sense is a non-trivial task.
+Our approach addresses this critical gap by providing a systematic framework for generating spatially coherent data.
 In this work, we introduce SKG2Data, a novel multimodal synthesis approach guided by spatial knowledge graphs, grounded in the concept of knowledge-to-data generation.
-SKG2Data automatically constructs a Spatial Knowledge Graph (SKG) to emulate human-like perception of spatial directions and distances, which is subsequently utilized to guide multimodal data synthesis.
-Extensive experiments demonstrate that data synthesized from diverse types of spatial knowledge, including direction and distance, not only enhance the spatial perception and reasoning abilities of MLLMs but also exhibit strong generalization capabilities.
-We hope that the idea of knowledge-based data synthesis can advance the development of spatial intelligence.
-Spatial Knowledge Graph-Guided Multimodal Synthesis
-Yida Xue1,  Zhen Bi1,2,  Jinnan Yang2,  Jungang Lou2, Huajun Chen1,  Ningyu Zhang1thanks:  Corresponding Author.
-1Zhejiang University  2Huzhou University {xueyida, bizhen_zju, huajunsir, zhangningyu}@zju.edu.cn
-1 Introduction
-Despite significant advancements in Multimodal Large Language Models (MLLMs) Zhao et al.
-(2023b); Yin et al.
-(2023) for visual processing tasks, a critical limitation persists in their ability to comprehend spatial relationships Prasad et al.
-(2023); Liu et al.
-(2023b); Du et al.
-(2024); Lei et al.
-(2024); Tong et al.
-(2024); Shiri et al.
-(2024); Zheng et al.
-(2024).
-Tong et al.
-(2024) observe that MLLMs often operate with “eyes wide shut”, revealing fundamental deficiencies in visual-spatial comprehension even in state-of-the-art models.
+SKG2Data employs an automated pipeline for constructing Spatial Knowledge Graph (SKG) that effectively captures human-like spatial cognition, including directional and distance relationships.
+These structured representations then serve as precise guidance for our integrated synthesis pipeline, where a diffusion model generates spatially-consistent images while a MLLM produces corresponding textual descriptions.
+The automated construction of SKG enables scalable generation of diverse yet realistic spatial configurations, overcoming the limitations of manual data collection and annotation.
+Extensive experiments demonstrate that data synthesized from diverse types of spatial knowledge, including direction and distance, enhance the spatial perception and reasoning abilities of MLLMs markedly, albeit with a slight cost to their general capabilities.
+We hope that the idea of knowledge-based data synthesis can advance the development of spatial intelligence [^1] Code is available at https://github.com/zjunlp/Knowledge2Data..
+I Introduction
+Despite significant advancements in Multimodal Large Language Models (MLLMs) [1, 2] for visual processing tasks, a critical limitation persists in their ability to comprehend spatial relationships[3, 4, 5, 6, 7, 8, 9].
+Recent studies demonstrate that even state-of-the-art MLLMs operate with “eyes wide shut” [7], exhibiting fundamental deficiencies in visual-spatial comprehension.
+These models show marked inabilities to accurately estimate relative positions between objects, consistently fail in judging spatial distances, and possess a limited capacity for reasoning about the spatial relationships among multiple objects.
 This limitation underscores a significant gap between MLLMs and human capabilities, as humans inherently excel in tasks requiring spatial intelligence.
@@
 To address this challenge, multimodal data synthesis offers a promising solution, requiring the generation of large-scale, spatially coherent multimodal data that accurately adheres to real-world constraints.
-A straightforward approach involves developing computational models (or world models) capable of generating continuous synthetic data Du et al.
-(2024); Ray et al.
-(2024); Li et al.
-(2025).
-However, this method is inefficient due to redundant data across scenarios and demands frequent model updates when environments change, leading to significant maintenance costs.
-Notably, human spatial intelligence stems from the brain’s acquisition of spatial knowledge, which enables robust generalization through the application of cognitive frameworks Hodges et al.
-(1999); Herweg and Kahana (2018); Peer et al.
-(2021).
+A straightforward approach involves developing computational models (or world models) capable of generating continuous synthetic data [5, 10, 11].
+However, this method proves inefficient due to redundant data across scenarios and the need for frequent model updates when environments change, resulting in substantial high computational costs for environment updates.
+In contrast, human spatial intelligence stems from the acquisition of spatial knowledge by the brain, which allows robust generalization through the application of cognitive frameworks [12, 13, 14].
 This insight naturally leads to a new paradigm for multimodal data synthesis: modeling spatial knowledge to guide the targeted generation of informative data.
-Inspired by this insight, as illustrated in Figure 1, we introduce SKG2Data, an innovative framework specifically designed to generate multimodal synthetic data for spatial understanding.
-The core idea is to leverage a specialized knowledge graph (KG) as a spatial knowledge representation to guide the generation of multimodal data.
-Specifically, we construct a Spatial KG (SKG) that simulates human-like understanding of spatial knowledge by characterizing scenes and modeling spatial relationships.
-SKG2Data utilizes the SKG, which encodes object attributes and their spatial relationships, to generate corresponding images and textual data.
-We automatically construct a multimodal instruction dataset for MLLMs based on the SKG, comprising 974 nodes and 4,731 triplets that encode spatial relationships.
+Inspired by this insight, we introduce SKG2Data, an innovative framework specifically designed to generate multimodal synthetic data for spatial understanding as illustrated in Figure 1.
+The core idea is to leverage a specialized knowledge graph as a structured spatial representation to guide multimodal data generation.
+Specifically, we construct a Spatial KG (SKG) by employing GPT-4o to characterize objects and model their spatial relationships, thereby establishing a structured representation of spatial knowledge.
+The resulting SKG, which encodes object attributes and their spatial relationships, then serves as the key control mechanism in SKG2Data to direct the generation of corresponding images and textual data.
+We automatically synthesize SKG-based spatial multimodal data and design three ways to ensure the quality of synthetic data: human verification, external knowledge base enhancement, and automatic multimodal agent verification.
+After filtering, we construct a multimodal instruction dataset, which contains 974 nodes and 4,731 triplets encoding spatial relationships.
 To ensure a fair comparison of results and prevent data contamination, we also use SKG2Data to synthesize an evaluation dataset, referred to as SKG2Data-Holdout, which includes 120 images and 566 single-choice questions.
-We fine-tune the LLaVA-1.6-Vicuna-7B Liu et al.
-(2024a) and Llama-3.2-Vision-11B Meta AI (2024) models using the synthesized data.
-The fine-tuned models show significant performance improvements on SKG2Data-Holdout and other spatial understanding benchmarks, while maintaining their general abilities.
+We fine-tune the LLaVA[15] and Llama-3.2-Vision[16] models using the synthesized data.
+The fine-tuned models show performance improvements on SKG2Data-Holdout and other spatial understanding benchmarks, while maintaining their general abilities.
 Additionally, we carry out comprehensive experiments to examine the effect of different spatial relation types, dataset sizes, and object densities on model performance.
+Our work makes the following principal contributions:
+  • We propose SKG2Data, a novel framework for generating multimodal synthetic data for spatial understanding tasks.
+It uses a SKG to guide the creation of high-quality multimodal data that reflects real-world spatial constraints, addressing the critical limitation of existing MLLMs in comprehending spatial relationships.
+  • We automatically construct a multimodal instruction dataset and an evaluation dataset SKG2Data-Holdout based on the SKG2Data.
+These datasets are rich and diverse, providing comprehensive training data and a fair benchmark for assessing the spatial understanding capabilities of MLLMs.
+  • We demonstrate the effectiveness of our synthesized data through extensive experiments.
+Fine-tuning MLLMs with our data significantly improves their performance on spatial understanding tasks while only slightly reducing their general capabilities.
+Our findings also reveal that directional knowledge is crucial for enhancing spatial perception and that increasing object density in the data improves fine-grained recognition.
+II Related Work
+II-A Multimodal Large Language Models
+The remarkable success of Large Language Models (LLMs) [17, 18, 19] has driven significant advances in Multimodal Large Language Models (MLLMs) [20, 15, 21, 16, 22, 23, 24, 25].
+These MLLMs typically employ transformer-based architectures with specialized visual encoders, such as ViT [26] or CLIP [27], that project visual inputs into the language model’s embedding space.
+This architectural innovation enables models to process and understand both text and visual information through a unified framework, achieving state-of-the-art performance on visual tasks.
+The development of comprehensive evaluation benchmarks has been crucial for measuring progress in this field.
+General-purpose multimodal benchmarks like MMVet [28] and MMStar [29] assess broad capabilities including visual recognition, text generation, and commonsense reasoning.
+More specialized benchmarks like HallusionBench [30] and POPE [31] focus on identifying and quantifying hallucinations in MLLMs.
+Recent work has highlighted spatial reasoning as a particularly challenging capability for MLLMs.
+This has led to the creation of dedicated spatial understanding benchmarks [32, 5, 7] that systematically evaluate abilities like relative position estimation, distance judgment, and multi-object spatial relationships.
+These benchmarks reveal that while MLLMs excel at many visual-language tasks, their spatial comprehension still lags significantly behind human performance, particularly in complex real-world scenarios requiring spatial reasoning.
+II-B Synthetic Data Generation
+Synthetic data has garnered significant attention as a novel solution to the challenges of obtaining large, diverse, and high-quality datasets [33, 34, 35, 36], particularly in domains where real-world data collection is expensive, privacy-sensitive, or inherently limited.
+Multimodal data synthesis represents an advanced and rapidly evolving field with transformative potential across computer vision, natural language processing, and human-computer interaction applications.
+This paradigm involves the systematic creation of integrated multimodal data, combining text, images, audio, and sensor modalities through algorithmic generation processes.
+Specifically, for current MLLMs, many works focus on synthesizing data in both image and text modalities to train or test models.
+Many existing studies primarily rely on pre-existing images [37, 38, 39], occasionally incorporating additional image information such as captions, bounding boxes, or OCR input [40, 41, 42].
+However, another area of research transcends generating textual instruction data and focuses on creating high-quality images through various synthesis techniques rather than relying solely on existing image datasets.
+For example, REACHQA [43] and Multimodal Self-Struct [44] use code to precisely synthesize chart images.
+Additionally, Scenethesis [45] and EmbSpatial [5] utilize indoor 3D simulators to generate images, while SynCLR [46] and VisMin [47] employ diffusion models to produce new images.
+Furthermore, since LLMs tend to exhibit biases and follow a long-tailed distribution due to their lack of specialization in specific domains [48, 49], there is a growing body of research [50, 51, 52, 53, 54] that employs knowledge enhancement techniques, such as KG, to generate higher quality data.
+In this work, we propose a method for multimodal data synthesis based on diffusion models and knowledge enhancement techniques.
+End-to-end text-to-image generative models, such as Stable Diffusion [55] and Dall-E [56], often encounter challenges in accurately controlling object positions and quantities.
+In contrast, layout-based diffusion models, such as GLIGEN-based models [57, 58], achieve accurate control over object placement and quantities by incorporating bounding boxes or explicit layout guidance.
+LLM-grounded Diffusion (LMD) [57] utilizes LLMs to generate bounding boxes as layout information for these layout-based diffusion models.
+Specifically, our method leverages spatial knowledge graphs to guide the generation of both images and corresponding textual descriptions, ensuring that the synthesized data adheres to real-world spatial constraints.
+Recent advances in diffusion models enable high-fidelity image generation, while knowledge graphs provide structured representations that maintain semantic and spatial coherence across modalities.
 Figure 2: A comprehensive overview of the our framework.
@@
 The Multimodal Data Synthesis module is tasked with generating image data and their corresponding textual data.
-2 Preliminaries
-Spatial Knowledge Graph.
-In recent years, KGs have been widely applied across various domains, such as FACTKG Kim et al.
-(2023) for fact verification.
+III Preliminaries
+III-A Spatial Knowledge Graph
+In recent years, KGs have been widely applied across various domains, such as FACTKG [59] for fact verification.
 Unlike traditional KGs, we introduce the Spatial KG (SKG), a specialized KG designed to encode spatial knowledge and guide multimodal data synthesis for spatial tasks.
@@
 SKG represents a collection of objects and their spatial relationships within a specific scene.
-In contrast to Scene Graphs Zhu et al.
-(2017, 2022), which provide a general representation of scenes, the triplets in an SKG are specifically designed to emphasize spatial relations, with nodes $E$ containing detailed descriptions of objects.
-This structured representation enables the generation of complex and detailed multimodal synthesis data that emphasize spatial relationships.
-Multimodal Data Synthesis.
+In contrast to Scene Graphs [60, 61], which provide a general representation of scenes, the triplets in an SKG are specifically designed to emphasize spatial relations, with nodes $E$ containing detailed descriptions of objects.
+This structured representation facilitates the generation of complex, detailed multimodal data that accurately capture spatial relationships.
+By systematically modeling spatial relationships through SKG, our method bridges the gap between multimodal synthetic data and human-like spatial understanding.
+III-B Multimodal Synthetic Data
 The task of multimodal data synthesis involves generating high-quality input-output pairs related to images using a model $f$.
@@
 Our methodology involves the simultaneous creation of images and their corresponding question-answer pairs through a three-component framework: a multimodal generative model $f_{1}$, a text-to-image generative model $f_{2}$, and a SKG $G$.
-Inspired by LLM-grounded Diffusion (LMD) Lian et al.
-(2024), we use $f_{1}$ to generate bounding boxes $B$ and description $d$ under the guidance of $G$.
+Inspired by LLM-grounded Diffusion (LMD) [57], we use $f_{1}$ to generate bounding boxes $B$ and description $d$ under the guidance of $G$.
 These bounding boxes are then integrated into the layout-diffusion model $f_{2}$ to synthesize images.
 After image synthesis, $G$ and $B$ further assist $f_{1}$ in generating multiple question-answer pairs $(Q,A)$ relevant to the synthesized image $i$.
+This process can be mathematically expressed as follows:
 $\begin{gathered}(B,d)=f_{1}(G)\\ i=f_{2}(B,d,G)\\ (Q,A)=f_{1}(B,G,i)\end{gathered}$ (2)
-For implementation, we use GPT-4o and GLIGEN Li et al. (2023) version of LMD as $f_{1}$ and $f_{2}$, respectively.
-3 Spatial Knowledge Graph-Guided Data Synthesis
+For implementation, we use GPT-4o and GLIGEN [58] version of LMD as $f_{1}$ and $f_{2}$, respectively.
+IV Spatial Knowledge Graph-Guided Data Synthesis
 In this section, we introduce SKG2Data, a novel approach that takes advantage of a SKG that incorporates spatial positional relationships to guide the synthesis of multimodal data.
-As shown in Figure 2, SKG2Data consists of two core modules: Spatial KG Generation and Multimodal Data Synthesis.
-3.1 Spatial Knowledge Graph Generation
-Scenes and Objects Generation.
+As shown in Figure 2, SKG2Data consists of two modules: Spatial KG Generation and Multimodal Data Synthesis.
+IV-A Spatial Knowledge Graph Generation
+Scenes and Objects Generation
 The real world encompasses a wide variety of scenes, each containing distinct sets of objects.
 To simulate the distribution of scenes and objects in real-world images, we leverage GPT-4o to generate a series of scenes and the objects likely to appear within them.
-Starting with a few scenes as few-shot examples, we use GPT-4o to generate an additional 160 safe, diverse, and realistic scenes, significantly enhancing the diversity of our dataset.
+Starting with a few scenes as few-shot examples, we use GPT-4o to generate an additional scenes.
+After human inspection and filtering, we finally obtain 160 safe, diverse, and realistic scenes, significantly enhancing the diversity of our dataset.
 For each scene, GPT-4o is further employed to produce a list of plausible objects that could exist in the given context.
-To improve object generation, particularly for uncommon scenes, we incorporate Wikipedia documents as external knowledge to augment GPT-4o’s understanding.
+To improve object generation, particularly for uncommon scenes, we incorporate Wikipedia documents as external knowledge to ensure the trustworthiness and rationality of objects in each scene.
 This approach ensures that the synthesized objects exhibit distributions closely aligned with real-world images, effectively preventing the generation of incongruous or improbable objects within the specified scenes.
+Through observation of the generated objects, we find that most of the objects appear in high frequency in each real scene.
 Formally, we first employ $f_{1}$ to generate $m$ general scenes, denoted as $\{s_{1},s_{2},\dots,s_{m}\}$.
@@
 $\{o_{i1},o_{i2},\dots,o_{in}\}=f_{1}(s_{i})$ (3)
-Figure 3: Distribution of top 15 objects and spatial relationships. There are a total of 974 objects and 95 relationships.
-Spatial Knowledge Graph Construction.
+Spatial Knowledge Graph Construction
 Due to the diverse rules and knowledge governing spatial relationships between objects in the real world, we employ GPT-4o to generate positional relationships that align with real-world spatial cognition.
@@
 $\begin{gathered}C_{k}=\{o_{ij_{1}},o_{ij_{2}},\ldots,o_{ij_{k}}\},\\ C_{k}\subseteq\{o_{i1},o_{i2},\ldots,o_{in}\},\\ |C_{k}|=k\end{gathered}$ (4)
+Figure 3: Distribution of top 15 objects and spatial relationships. There are a total of 974 objects and 95 relationships.
+Figure 4: Training data cases featuring diverse spatial relationships, synthesized through our proposed method.
 Next, each object in $C_{k}$ undergoes an attribute enrichment process via $f_{1}$, which adds quantities and attributes such as color, orientation, and material.
@@
 By combining the nodes $E_{l}$ and the triplet relationships $T_{l}$, we construct the $G$.
-3.2 Multimodal Data Synthesis
-Image Data Generation.
-End-to-end text-to-image generative models, such as Stable Diffusion Rombach et al.
-(2021) and Dall-E Betker et al.
-(2023), often encounter challenges in accurately controlling object positions and quantities.
-In contrast, layout-based diffusion models enable precise control over object placement and quantities, ensuring that the generated images closely conform to SKG.
+IV-B Multimodal Data Synthesis
+Image Data Generation
 We employ GPT-4o to generate bounding boxes and captions under the guidance of SKG, and seamlessly incorporate these outputs into a GLIGEN-based model.
@@
 Additionally, we leverage few-shot examples to instruct $f_{1}$ on how to generate appropriate bounding boxes $B_{l}$.
-$\begin{gathered}B_{l}=\{b_{ij_{1}},b_{ij_{2}},\ldots,b_{ij_{l}}\}=f_{1}(E_{l},T_{l}),\\ \end{gathered}$ (7)
+[…]
 In order to synthesize data on whether an object exists in the image for subsequent tasks, we randomly select an object from other objects in current scene as a non-existent object $o_{neg}$, ensuring that this object is excluded from the synthesized image.
 The elements $E_{l}$, $T_{l}$, and $s_{i}$, along with the optional non-existent object $o_{neg}$, are input to $f_{1}$ to produce the caption $d$.
+The formula is as follows:
 $\begin{gathered}o_{neg}\in\{o\mid o\notin C_{k}\}\cup\{\emptyset\},\\ d=f_{1}(E_{l},T_{l},s_{i},o_{neg})\end{gathered}$ (8)
@@
 $\begin{gathered}i=f_{2}(d,B_{k})\end{gathered}$ (9)
-Image Data Filtering.
+Image Data Filtering
 During the image generation stage, a significant challenge is the potential for discrepancies, such as hallucinations, which can occur during the transformation of the SKG into bounding boxes and captions.
@@
 This process is aimed at evaluating and ensuring that the generated image aligns with the SKG.
-Formally, we utilize $f_{1}$ to assess the alignment of an image $i$ with the expected entities $E_{l}$ and textual descriptions $T_{l}$.
+Formally, we utilize $f_{1}$ to assess the alignment of an image $i$ with the expected entities $E_{l}$ and relationships $T_{l}$.
 If $f_{1}$ returns True, the image $i$ is preserved; otherwise, it is discarded.
@@
 $f_{1}(i,E_{l},T_{l})=\begin{cases}\text{True},&\text{if }i\text{ aligns with}\\ &\text{both }E_{l}\text{ and }T_{l},\\ \text{False},&\text{otherwise.}\end{cases}$ (10)
+By verifying the attributes and positional relationships of objects in the image through the agent, filter out images that do not match the information in SKG.
 After filtering, the dataset contains a total of 974 objects and 95 spatial relationships.
 The distribution of the top 15 objects and spatial relationships in the filtered images is illustrated in Figure 3.
+TABLE I: Comparison of different models on five benchmarks.
 Dataset LLaVA-1.5 LLaVA-1.6 LLaVA-1.6 (w/ SKG2Data) Llama-3.2-vision Llama-3.2-vision (w/ SKG2Data) KG2Data-Holdout 62.2 68.6 70.1 (+1.5) 73.3 74.7 (+1.4) COCO-Spatial 52.4 75.4 79.3 (+3.9) 45.9 59.8 (+13.9) MMVP 24.7 32.0 36.7 (+4.7) 29.3 30.7 (+1.4) MMStar 33.1 37.6 36.7 (-0.9) 49.8 48.1 (-1.7) HallusionBench 27.6 27.6 27.2 (-0.4) 40.3 45.1 (+4.8) Average 40.0 48.2 50.0 (+1.8) 47.7 51.7 (+4.0)
-Table 1: Comparison with SoTA methods on five benchmarks.
-Textual Data Generation.
+Examples of the training data generated are illustrated in Figure 4.
+Textual Data Generation
 The training data for textual data is primarily composed of open-ended question-answer pairs, while the evaluation data predominantly consists of single-choice questions.
@@
 These questions are categorized into two types based on their primary focus: Entity-Based Data and Relation-Based Data.
-[…]
+Entity-Based Data focuses on the existence, attributes, and quantities of objects in a given image.
+The construction of Entity-Based Data is primarily guided by the entity set $E_{l}$, along with explicitly defined non-existent objects $o_{neg}$.
+This type of data serves as general visual question-answering data, maintaining the general capabilities of MLLMs without requiring additional external datasets.
+In contrast, Relation-Based Data emphasizes the spatial relationships between objects.
+The generation of Relation-Based Data is based on the spatial relationship triplets $T_{l}$ and generated bounding boxes $B_{l}$.
+Relation-Based Data is specifically designed to improve the spatial reasoning capabilities of MLLMs.
+The generation process for each question $x$ and its corresponding answer $y$ is formalized as follows:
 $\begin{gathered}x=f_{1}(i,B_{l},E_{l},T_{l}),\\ y=f_{2}(i,B_{l},E_{l},T_{l},x)\end{gathered}$ (11)
-Textual Data Filtering.
+Textual Data Filtering
 For each generated textual data, we employ GPT-4o as a text filtering agent to verify its accuracy.
@@
 $f_{1}(i,x,y)=\begin{cases}\text{True}&\text{if }y\text{ is correct},\\ \text{False}&\text{otherwise.}\end{cases}$ (12)
-Examples of the final training data generated are illustrated in Figure 7 (Appendix A).
-4 Experiments
-4.1 Experimental Settings
-Baselines.
-We conduct experiments on three open-source MLLMs: LLaVA-1.5-7B Liu et al.
-(2023a), LLaVA-1.6-7B Liu et al.
-(2024a), and Llama-3.2-Vision-11B Meta AI (2024).
-Both LLaVA-1.5 and LLaVA-1.6 use the Vicuna-7B Chiang et al.
-(2023) LLM as their LLM backbone.
-Benchmarks.
+After filtering, data pairs that the GPT-4o cannot correctly answer solely by providing images can be filtered.
+Figure 5: The result of removing specific positional relationship data with the same amount of data.
+V Experiments
+V-A Experimental Settings
+Baselines
+We conduct experiments on three open-source MLLMs: LLaVA-1.5-7B [40], LLaVA-1.6-7B [15], and Llama-3.2-Vision-11B [16]. Both LLaVA-1.5 and LLaVA-1.6 use the Vicuna-7B [62] LLM as their LLM backbone.
+Benchmarks
 For a comprehensive evaluation, we examine all models across two distinct task categories.
 The first category encompasses benchmarks that are either directly or partially related to spatial understanding.
-This includes the publicly available benchmark MMVP and COCO-Spatial Kamath et al.
-(2023).
+This includes the publicly available benchmark MMVP [7] and COCO-Spatial [32].
 For COCO-Spatial, we utilize a subset named Two-obj that focuses on the positional relationships of two objects.
@@
 We leave a portion of the synthesized image to generate test questions called SKG2Data-Holdout, which are also used for evaluation.
-The second category involves general capability testing benchmarks for MLLMs, including MMStar and HallusionBench.
-Implementation Details.
+The second category involves general capability testing benchmarks for MLLMs, including MMStar [29] and HallusionBench [30].
+Implementation Details
 We fine-tune both LLaVA-v1.6 and Llama-3.2-Vision using training data.
 These models consist of a vision encoder, a projector, and an LLM backbone.
-During training, we tune the LLM backbone with LoRA modules  Hu et al.
-(2021), keeping the projector and vision encoder parameters fixed.
-The detailed settings are shown in Table 2 (Appendix A).
-4.2 Main Results
-Table 1 presents a comprehensive comparison of multiple models evaluated across five benchmarks, highlighting the strengths of SKG2Data.
-Spatial Understanding Results.
+During training, we tune the LLM backbone with LoRA modules [63], keeping the projector and vision encoder parameters fixed.
+The experimental setup employs LoRA for fine-tuning with a global batch size of 64.
+Training occurs on NVIDIA A800 GPUs using a learning rate of $1\times 10^{-5}$.
+The LoRA configuration includes a rank ($r$) of 16, an alpha ($\alpha$) value of 32, and a dropout rate of 0.05.
+The models undergo training for 2 epochs with mixed precision enabled to optimize computational efficiency.
+These parameters ensure stable training while maintaining memory efficiency and model performance.
+Figure 6: The influence of the number of objects and the quantity of data on experimental results.
+V-B Main Results
+Table I presents a comprehensive comparison of multiple models evaluated across five benchmarks, highlighting the strengths of SKG2Data.
+Spatial Understanding Results
 Fine-tuned LLaVA-1.6 and fine-tuned Llama-3.2-vision demonstrate marked improvements across several challenging benchmarks for spatial tasks, highlighting the efficacy of our methods.
@@
 In the MMVP benchmark, Fine-tuned LLaVA-1.6 shows notable improvement with a score of 36.7(+4.7)↑, while Fine-tuned Llama-3.2-vision also makes gains, reaching 30.7(+1.4)↑.
-In summary, our experimental results indicate that synthesized training data can improve the spatial understanding ability of MLLMs.
+In summary, our experiments show that synthesized training data enhances MLLMs’ spatial understanding, even without scene or object generation tailored for these spatial datasets.
 General Visual Understanding Results.
@@
 Average Results.
-The fine-tuned LLaVA-1.6 model achieves an average score of 50.0 (+1.8), while the fine-tuned Llama-3.2-vision model reaches 51.7 (+4.0).
+Many current works have found that training on datasets with specific domains or abilities can affect the performance of models in other domains and general domains, leading to an overall performance declines [64, 65, 66].
+For our method, the fine-tuned LLaVA-1.6 model achieves an average score of 50.0 (+1.8), while the fine-tuned Llama-3.2-vision model reaches 51.7 (+4.0).
 These improvements confirm that SKG2Data effectively enhances spatial reasoning while maintaining performance in general visual understanding tasks.
-Figure 4: The result of removing specific positional relationship data with the same amount of data.
-Figure 5: The influence of the number of objects and the quantity of data on experimental results.
-4.3 Ablation and Analysis
+Figure 7: Cases analysis: In spatially related datasets, the trained model enhances spatial understanding abilities.
+V-C Ablation and Analysis
 To evaluate the impact of the synthesized data generated by SKG2Data on model performance, we conducted a comprehensive ablation study using LLaVA-1.6.
@@
 Additionally, Finding 3 is derived from the experimental results observed in Figure 6.
-Figure 6: Case analysis: On spatially related datasets, the trained model enhances spatial understanding abilities.
-Finding 1: Directional knowledge plays a crucial role in enhancing the MLLM’s spatial perception capabilities.
+Finding 1. Directional knowledge plays a crucial role in enhancing the MLLM’s spatial perception capabilities
 We observe that training data containing directional relationships leads to more significant performance improvements compared to data with distance relationships across three spatially correlated datasets.
@@
 A potential explanation is that directional knowledge, compared to distance knowledge, exhibits greater variability, providing stronger spatial signals that are more effective in enhancing model performance.
-Finding 2: Data synthesized from two types of spatial knowledge, including direction and distance, exhibit generalization ability.
+Finding 2. Data synthesized from two types of spatial knowledge, including direction and distance, exhibit generalization ability
 It can be observed that the use of two types of spatial relationship data can enhance performance across three datasets.
@@
 This phenomenon underscores the superiority of knowledge-based data synthesis over previous methods.
-Finding 3: Increasing the number of objects improves the performance of MLLMs in visual detail tasks
+Finding 3. Increasing the number of objects improves the performance of MLLMs in visual detail tasks
 Our experiments demonstrate that synthesizing data with varying quantities of objects improves spatial understanding, with greater gains observed as the number of objects increases in MMVP benchmark.
@@
 By leveraging the structured spatial knowledge embedded in SKG, our proposed SKG2Data systematically generates multimodal training data with diverse and spatially plausible object arrangements, enabling more effective learning of visual and spatial reasoning.
-Discussion on Knowledge-to-Data Synthesis.
-It can be seen that with more data, the performance of the trained model can be further improved.
-At the same time, it can also be found that high-quality data is important, and a small amount of high-quality training data can significantly improve the performance of models in specific fields, which is in line with the opinion of “less is more” Zhou et al.
-(2023); Ye et al.
-(2025).
-As shown in Figure 6, the trained model exhibits improved accuracy on spatially related questions.
+Noise and Error Analysis
+Despite the implementation of automatic filtering, the synthesized data may still contain noise.
+This noise arises from potential mismatches between images and text in diffusion models, as well as the inherent possibility of hallucinations in MLLMs.
+In addition, due to the inherent limitations of current diffusion models, generated images often show problems like distortion and inconsistencies in object type, attributes, quantity, or spatial arrangement.
+Creating realistic images becomes much harder as the SKG becomes more complex.
+Specifically, when the SKG grows larger—with more nodes representing objects and more edges encoding their relationships—the image generation task becomes significantly more difficult.
+Therefore, the various SKG graph structures currently sampled for synthesizing training data are not too complex.
+In some real-world situations where scenes are packed with many objects and complex spatial relationships, our method may have difficulty accurately generating these complex data.
+Figure 8: The mechanism of changes in the LLaVA model before and after training.
+We use two cases to calculate the probability of tokens representing various options at each layer of the transformer, which reveals that the trained model improves the probability of correct options in the last few layers, especially in the last layer.
+Discussion on Knowledge-to-Data Synthesis
+As shown in our experiments, increasing the volume of training data generally leads to improved model performance.
+Moreover, we observe that high-quality data is important and a small amount of high-quality training data can significantly improve the performance of models in specific fields, which is in line with the opinion of “less is more” [67, 68].
+As illustrated in the three cases of Figure 7, the trained model demonstrates significantly improved accuracy in spatially related datasets.
+To further investigate how the LLaVA model evolves during training, we select two types of case: one that the model answers correctly even before training and another that it answers correctly only after training.
+We analyze internal representations [69, 70, 71] in different layers by processing these inputs through the model.
+An LM-head attached to each layer produces logits for the first generated token, which are then converted into probabilities via softmax.
+As shown in Figure 8, the token probability distribution remains largely unchanged in the previous transformer layers.
+However, in the last few transformer layers, especially in the last layer, the probability of correct token is rapidly improved.
 The core idea of our work is to synthesize data through automated modeling knowledge to ensure the rationality of the data itself.
@@
 In addition, to synthesize additional data, we can either randomly sample more data from the SKG or expand the scenes to further enrich the SKG.
-Compared with the method of synthesizing multimodal data using code, simulators, etc., it is more efficient and has better generalization ability.
-5 Related Work
-Multi-Modal Large Language Model.
-The remarkable success of large language models (LLMs) Brown et al.
-(2020); Achiam et al.
-(2023); Touvron et al.
-(2023) has driven the development of MLLMs Ye et al.
-(2023); Liu et al.
-(2024a); Chen et al.
-(2023); Meta AI (2024); Gemini Team (2024); Bai et al.
-(2023); OpenAI (2023, 2024).
-By incorporating visual backbones, these MLLMs demonstrate strong performance across a wide range of visual-language tasks, significantly improving the processing and understanding of both text and visual information.
-At the same time, a variety of benchmarks Yu et al.
-(2023); Guan et al.
-(2024); Chen et al.
-(2024) have been developed to evaluate the capabilities of large multimodal models.
-Additionally, with increasing attention on spatial understanding in MLLMs, specialized benchmarks Kamath et al.
-(2023); Du et al.
-(2024); Tong et al.
-(2024) have been introduced.
-These benchmarks are specifically designed to tackle the challenges of evaluating spatial comprehension in MLLMs.
-Synthetic Data Generation.
-Synthetic data has garnered significant attention as a novel solution to the challenges of obtaining large, diverse, and high-quality datasets Chang et al.
-(2024); Liu et al.
-(2024b); Long et al.
-(2024); Kim et al.
-(2024).
-Many existing studies primarily rely on pre-existing images Wu et al.
-(2024), occasionally incorporating additional image information such as captions, bounding boxes, or OCR input Liu et al.
-(2023a); Zhao et al.
-(2023a); Zhang et al.
-(2023).
-However, another area of research transcends generating textual instruction data and focuses on creating high-quality images through various synthesis techniques rather than relying solely on existing image datasets.
-For example, REACHQA He et al.
-(2024) and Multimodal Self-Struct Zhang et al.
-(2024) use code to precisely synthesize chart images.
-Additionally, EmbSpatial utilizes simulators to generate images, while SynCLR Tian et al.
-(2024) and VisMin Awal et al.
-(2024) employ diffusion models to produce new images.
-Furthermore, due to the tendency of LLMs to exhibit biases and follow a long-tailed distribution due to their lack of specialization in specific domains Navigli et al.
-(2023); Fei et al.
-(2023), there is a growing body of research Feng et al.
-(2023); Xu et al.
-(2024); Wang et al.
-(2024); Yang et al.
-(2025); Ovadia et al.
-(2025) that employs knowledge enhancement techniques, such as KG, to generate higher-quality data.
-6 Conclusion
+Most existing simulators [72, 73, 74, 75, 76, 77, 78] are designed for specific environments, and image synthesis via code is largely confined to mathematical visualizations or web-based content.
+Consequently, such approaches are unable to effectively generate multimodal spatial data beyond these limited contexts.
+In contrast, our method supports effective data generation across a wide variety of scenarios.
+VI Conclusion
 In this work, we propose SKG2Data, a novel multimodal synthetic data generation method that leverages Spatial Knowledge Graph to synthesize images and question-answer pairs, addressing the lack of high-quality data for spatial understanding tasks.
 Using SKG2Data, we create a multimodal instruction dataset and a benchmark, SKG2Data-Holdout, to enhance and evaluate the spatial understanding capabilities of MLLMs.
-Extensive experiments demonstrate the effectiveness of SKG2Data and analyze the impact of spatial relationship types, dataset scales, and object densities.
-Limitations
-Noise in synthetic data.
-Despite the implementation of automatic filtering, the synthesized data may still contain noise.
-This noise arises from potential mismatches between images and text in diffusion models, as well as the inherent possibility of hallucinations in LLMs.
-Difficult to synthesize data with a large number of objects.
-Given the limitations of current diffusion models, generated images can sometimes show problems like distortion and inconsistencies in object type, attributes, quantity, or spatial arrangement.
-The challenge of creating realistic images grows substantially as the number of objects in the scene increases.
-In some real-world situations where scenes are packed with many objects, our method may have difficulty accurately depicting these complex scenes.
-Appendix A Appendix
-Hyperparameter Setting
-Fine-tuning method LoRA
-Gobal Batch Size 64
-[…]
-Learning Rate (LR) $51e-$ 5
-LoRA $r$ 16
-LoRA $\alpha$ 32
-LoRA Dropout 0.05
-Epoch 2
-Mixed precision Enabled
-Table 2: Detailed experimental settings.
-Figure 7: Training data cases.
-Prompt for "Scenes Generation":
-You are an AI assistant specializing in imagining the scenes depicted in photos.
-Your task is to envision a variety of scenes that may appear in different photos.
-Your response needs to meet the following requirements: 1.
-Ensure that every scene is safe, accessible to most people, and common in real and daily life scenes.
-2.
-The scenes in the photos should be diverse and non repetitive.
-Each generated scene typically depicts a facility (usually requiring interaction with multiple objects such as a table or furniture) or a room, or a place or area.
-Please directly output the generated scenes in the following list format: <In context learning> Now, please generate 160 different scenes.
-<image>
-Prompt for "Objects Generation":
-<Wikipedia documents> You are an AI assistant specializing in identifying objects likely to appear in a given scene.
-Your task is to list as many objects as possible that may be found in that scene.
-Your response needs to meet the following requirements: 1.
-Ensure that each object is based on common situations in reality and daily life.
-Ensure that objects are relatively simple and easy to draw, and once you see their images, you can determine what objects they are.
-Ensure that the object is safe and harmless.
-2.
-The objects provided should not be duplicated or too similar.
-Try to describe an object with one word.
-If there are words with multiple meanings, it is necessary to ensure that they are distinguishable, such as "animal mouse" and "computer mouse".
-3.
-Each scene needs to list as many objects as possible.
-If there may be people on site, please add ’person’.
-Here I will give you some examples for reference and use the following template in the format of a list for the response: <In context learning> <image>
-Prompt for "Spatial KG Generation (Nodes)":
-You are an AI assistant specializing in describing objects in a given scene.
-Your task is to first select a subset of candidate objects, and then describe the attributes of the selected objects.
-Your response should meet the following requirements: 1.
-Select 2 to 4 suitable objects from the given objects to form a possible combination.
-Note that for some candidate objects that are suitable for repeated appearance, they can be selected repeatedly.
-Please prioritize selecting more common objects.
-Do not choose objects that are difficult to represent or recognize in the image.
-2.
-Provide further attribute descriptions for the selected objects, including color, orientation, etc., and describe countable objects as singular.
-Be careful not to describe objects that do not exist in the real world or are unsuitable for the current scene.
-The description of each object should be concise, without being overly complex and abstract, and avoid including details that cannot be visualized in the image.
-For some objects that are difficult to add descriptions, additional descriptions can be omitted.
-3.
-The selected combination of objects can be repeated, but the described attributes need to be different.
-4.
-If there are words with multiple meanings, it is necessary to ensure that they are distinguishable and you need to determine which one is more likely based on the scene and other objects, such as ’animal mouse’ and ’computer mouse’.
-5.
-You need to provide 25 to 50 possible combinations based on the number of objects given.
-Here I will give you some demonstrations for reference and use the following list template for the response: <In context learning> <image>
-Prompt for "Spatial KG Generation (Triplets)":
-You are an AI assistant specializing in describing the positional relationship between objects in a scene.
-Your task is to describe the positional relationship of the given objects.
-Your response needs to meet the following requirements: 1.
-Be careful not to describe positional relationships that do not exist in the real world.
-When describing the positional relationship between two objects, do not involve other objects.
-At most one positional relationship triplet is generated between every two objects.
-2.
-There may be two types of positional relationships between two objects: directional relationships and distance relationships.
-Directional relationships include above, below, left, right, in front of, behind, etc.
-Distance relationships include close, far, near, etc.
-Separate two positional relationships with commas when generating them simultaneously.
-For situations with a small number of objects, prioritize generating directional relationships.
-Here I will give you some demonstrations for reference and use the following template in the format of a list for the response: <In context learning> <image>
-Table 3: The prompt examples that we use in the first module.
-Prompt for "Image Data Generation":
-You are an intelligent bounding box generator.
-I will provide you with the scene, objects and their spatial relationships in a photo.
-Your task is to generate the bounding boxes for the given objects, along with a background prompt describing the scene and a prompt describing the photo.
-The images are of size 512x512.
-The top-left corner has coordinate [0, 0].
-The bottom-right corner has coordinnate [512, 512].
-The bounding boxes should not overlap or go beyond the image boundaries.
-Each bounding box should be in the format of (object name, [top-left x coordinate, top-left y coordinate, box width, box height]) and should not include more than one object.
-Do not put objects that are already provided in the bounding boxes into the background prompt.
-Do not include non-existing or excluded objects in the background prompt.
-Use "A realistic scene" as the background prompt if no background is given in the prompt.
-There may be multiple spatial relationships between objects, and special attention should be paid to using bounding boxes to represent these spatial relationshipss.
-Please refer to the example below for the desired format.
-<In context learning> <image>
-Prompt for "Image Data Filtering and Textual Data Generation":
-You are an AI assistant specializing in evaluating the quality of a given image and generating question and answer pairs related to the image.
-Your task is to first determine whether the given image should be saved based on its quality and whether it matches the descriptions, and then continue generating question and answer pairs if it is determined to be saved.
-Descriptions of the image will provide information about the objects present in the image, their quantity, their attributes, their positional relationships, their bounding boxes (object name, [top-left x coordinate, top-left y coordinate, box width, box height]), and objects that are absent from the image.
-You need to follow the steps below to complete the task: Step 1.
-You need to ensure the clarity, quality, and authenticity of the image.
-You need to verify if there are any unrealistic or unreasonable parts in this image, such as clearly unrealistic looking humans or animals, and if there are any indistinguishable objects.
-Step 2.
-You need to ensure that the image contains the specified objects and does not contain any excluded objects.
-Specified objects are given in the form of dict, and excluded objects are presented in a list format.
-Please note that the image should accurately reflect the attributes of the specified object, such as Orientation and Direction, State and Condition, Color and Appearance, Structural and Physical Characteristics, Viewpoint and Perspective.
-Step 3.
-You need to ensure that the objects in the image have a positional relationship with the matching descriptions.
-The positional relationships are given in the form of a list, and the objects correspond to the dict.
-Step 4.
-If the image matches step 1 and basically matches steps 2 and 3, please answer ’Yes’ and proceed to the next steps to generate question answer pairs.
-Otherwise, please answer ’No’ and do not proceed to the next steps.
-Step 5.
-You need to generate as many question answer pairs as possible based on the correct parts of these descriptions and information that can be confidently determined in the given image, while ensuring quality and diversity.
-For Step 5, the following requirements must be met: 1.
-Please note not to mention any information related to the image descriptions in your question-answer pairs, as the image description will not be provided in future tasks using these question-answer pairs.
-4.
-Use the template in the format of a JSON list to generate question and answer pairs without any other information.
-Below are reference responses for saving and not saving this image: <In context learning> <image>
-Prompt for "Textual Data Filtering":
-You are an AI assistant specializing in reviewing question-answer pairs about a given image.
-Your task is to find the correct question-answer pairs based on the given image.
-Your response needs to meet the following requirements: 1.
-You need to determine whether the answers to the questions are correct based on the given images, and save the correct question-answer pairs.
-2.
-Use a JSON list format template with the same format as the input to generate correct question-answer pairs without generating any additional information.
-Here is the question and answer about the given image: <In context learning> <image>
-Table 4: The prompt examples that we use in the second module.
+Extensive experiments analyze the effectiveness of SKG2Data and the impact of spatial relationship types, dataset scales, and object densities.
+Our results show that SKG2Data significantly improves MLLMs’ performance on spatial reasoning tasks.
+Furthermore, the automatic construction of Spatial Knowledge Graphs and data synthesis pipeline offers a flexible framework that can be extended to various domains requiring precise spatial representations.
+Acknowledgments
+We would like to express gratitude to the anonymous reviewers for their hard work and kind comments.
+This work was supported by the National Natural Science Foundation of China (No.
+NSFCU23B2055, No.
+62576307, No.62506128), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China under Grant (No.
+LRG25F030003 and LQN25F020023), Yongjiang Talent Introduction Programme (2021A-156-G), Ningbo Natural Science Foundation (2024J020), Information Technology Center and State Key Lab of CAD&CG.